OpenStudy (anonymous):

Put (7,-4) and (4,8) into Point Slope Form Please Help

OpenStudy (jdoe0001):

point-slope form => \(\huge {y-y_1=m(x-x_1)\\ (x_1,y_1), (x_2, y_2)}\)

OpenStudy (anonymous):

find the slope first slope(m) = (y2 - y1) / (x2 - x1) (7,-4) -- x1 = 7, y1 = -4 (4,8) -- x2 = 4, y2 = 8 sub all that in the formula. Once you find the slope(m), then sub the slope and either point (x1,y1) into this formula : y - y1 = m(x - x1), then you will have your answer. any problems, tag me and I will be happy to help :)